[figure: 9.7 cm]

SUBSTITUTION STUDY ON THE COHERENT KONDO STATE

IN CeCu_{5}In

D. Kaczorowski^{a}, Ya. M. Kalychak^{b} and V. I. Zaremba^{b} aW. Trzebiatowski Institute of Low Temperature and Structure Research, Polish Academy of Sciences, P. O. Box 1410, 50-950 Wrocław, Poland

bDepartment of Inorganic Chemistry, Lviv State University, 290 005 Lviv, Ukraine

The compound CeCu_{5}In crystallises in an ordered orthorhombic structure of its own type, which is closely related to that of CeCu_{6}. The magnetic, electrical transport and specific heat properties of this ternary phase indicate the formation at low temperatures of a coherent Kondo state without long-range magnetic order down to 1.7 K. Upon substitution of In or Ni for Cu the crystal structure type does not change. However, in CeCu_{5−x}In_{1+x} alloys two crystallographic sites have mixed Cu/In occupation, whereas in CeCu_{5−x}Ni_{x}In alloys probably all the transition metal sites are occupied by a statistical mixture Cu/Ni. These structural features have an enormous effect on the physical behaviour in the two systems. While the coherent Kondo state is destroyed immediately upon even very small Cu→In exchange, it survives the Cu→Ni substitution up to the Ni-atom concentration x = 0.4. Interestingly, all the alloys studied remain paramagnetic down to the lowest temperatures and exhibit a pronounced Kondo effect in the whole temperature range. The results obtained here for CeCu_{5−x}In_{1+x} and CeCu_{5−x}Ni_{x}In will be compared to those found previously for magnetically diluted alloys Ce_{1−x}La_{x}Cu_{5}In as well as for substitutional alloys CeCu_{5}In_{1−x}Ga_{x} and CeCu_{5}In_{1−x}Al_{x}.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : D. Kaczorowski

Address for correspondence : W.Trzebiatowski Institute of Low Temperature and Structure Research, Polish Academy of Sciences, P.O.Box 1410, 50-950 Wrocław, Poland

Email address : dkaczor@int.pan.wroc.pl